# Addressing the Non-perturbative Regime of the Quantum Anharmonic Oscillator by Physics-Informed Neural Networks

### v1

@@ III.3 Choice of the domain
We need to choose an $L$ such that the wave function is vanishingly small at the boundaries, but at the same time, $L$ should be as small as possible to avoid training the network on areas of just 00s, which would make the network vastly underestimate the orthogonality loss (9) when using Mean Squared Error. This means that the optimal choice of $L$ depends on the state that we want to train the network to predict, since for this type of potential higher energy states will tend to be less localized and we must therefore choose a domain that grows as $n$ increases [^3] If this means that we need to extrapolate, for instance when obtaining the values of $\psi_{i}(x)$ for the orthogonality loss, we just predict 0 outside the original computational domain since the wave function should be negligible outside it.. In particular, we empirically find that a good value for $L$ for the unperturbed oscillator is:
@@
To that we add a constant $\epsilon=1$ on each side of the interval to ensure some leeway [^4] But if this makes $L_{a}(n,\lambda)>L(n)$, we just use $L(n)$, thus obtaining:
@@ IV Results for the harmonic oscillator
We train the network to predict the first six stationary states. The training is stopped as soon as the total loss becomes less than $0.1$ and the differential equation loss is smaller than $3\times 10^{-3}$, in order to obtain good guesses for the energies and eigenstates without requiring an overly long computational time. We set $E_{init}=0$ and $a=0.8$ in (11), and obtain the values reported below in Table 2:
@@
0 $7.45\times 10^{-4}\pm\times 10^{-5}$ $0.9999926\pm 2.38\times 10^{-10}$
1 $-2.79\times 10^{-3}\pm 8.18\times 10^{-6}$ $0.9999751\pm 3.20\times 10^{-10}$
2 $-4.87\times 10^{-3}\pm 1.25\times 10^{-5}$ $0.9998467\pm 2.64\times 10^{-9}$
3 $-6.70\times 10^{-3}\pm 9.79\times 10^{-6}$ $0.9997815\pm 3.61\times 10^{-9}$
4 $-6.05\times 10^{-3}\pm 7.21\times 10^{-6}$ $0.9997674\pm 4.19\times 10^{-9}$
5 $-8.30\times 10^{-3}\pm 1.16\times 10^{-5}$ $0.9997969\pm 3.79\times 10^{-9}$
@@
Figure 1: Final wave function for the ground state. The blue line is the analytic wave function, while the red dots are the neural network prediction. In the left inset (in blue): the behavior of the loss throughout training, in logarithmic scale. In the right inset (in purple): fidelity throughout training
@@
Figure 2: Final wave function for the fifth excited state. The blue line is the analytic wave function, while the red dots are the neural network prediction. In the inset: the behavior of the loss throughout training, in logarithmic scale. In the right inset (in purple): fidelity throughout training
@@
The error on the eigenenergy always stays beneath the 1%. The results for the fidelity are also more than acceptable, with a distance of at most in the order of $10^{-4}$, even though the fidelity slightly decreases as the energy increases. Some examples of the final results and the behavior of the loss and fidelity throughout the training process are reported in Figure 1 for the ground state, and in Figure 2 for the fifth excited state. In particular, looking at the behavior of the loss we can see it is mostly monotonically decreasing, but for $n=5$ it takes roughly double the epochs then $n=0$ to get to the same value. Still, it shows good convergence. Concerning fidelity, it displays excellent behavior for $n=5$, asymptotically approaching 1. On the other hand, the behavior for $n=0$ is more complex. It rapidly approaches $0.8$, then goes back down to almost 00, and eventually rises up until it reaches 1. A possible explanation could be that the network was approaching the ground state, then it risked converging to an excited state, which will be orthogonal to the ground and therefore with fidelity zero, but after several epochs it managed to converge to the correct state, thanks to the energy minimization condition.
@@ V.1 Oscillator with real frequency
We now add the quartic potential to the oscillator. In particular, we will use $\omega=1$ and train the neural network with $\lambda=0.005\times 2^{n}$, with $n$ from 00 to $12$. These choices of $\lambda$ will allow us to study cases in which the quartic potential is just a perturbation, in which the two potentials have the same scales and in which the quartic part is dominant, up until the point where the system is basically just a pure quartic oscillator. In order to guarantee convergence, we exploit the results of section (IV) and employ transfer learning: we initialize the weight of the network to the ones that give the eigenfunction and eigenvalue of the harmonic oscillator, thus ensuring that the network starts close to the minima. Then, as $\lambda$ increases, we always initialize the network to the weight and biases as the already solved one with the closest $\lambda$. Due to this initialization, it becomes less important to enforce the orthogonality and energy minimization losses, since the minima closer to the starting state should already be the one corresponding to the correct eigenstate and eigenenergy, which means that we do not need these two losses as much to drive the network away from unwanted minima. Therefore, their starting weights have been decreased to $100n$ and $0.1$ respectively. The network manages to reach the convergence conditions for all values of $\lambda$ and $n$. Thanks to transfer learning, these results were also reached much quicker (5 to 10 minutes for lower excited states, compared to 20 to 40 minutes when given a clean initialization).
@@
0 0.005 $2.06\times 10^{-2}\pm 2.47\times 10^{-5}$
5 0.005 $-8.67\times 10^{-4}\pm 5.18\times 10^{-7}$
2 0.16 $1.56\times 10^{-2}\pm 1.27\times 10^{-5}$
2 0.32 $-1.87\times 10^{-3}\pm 8.34\times 10^{-6}$
2 0.64 $-4.79\times 10^{-2}\pm 8.04\times 10^{-6}$
3 1.28 $-2.41\times 10^{-1}\pm 3.16\times 10^{-6}$
0 10.24 $-4.18\times 10^{-1}\pm 1.03\times 10^{-5}$
5 10.24 $-7.94\times 10^{-1}\pm 6.21\times 10^{-6}$
@@
At this point we compare the obtained eigenenergies with those one gets from the standard perturbation theory. In particular, we will have, for $\omega=1$ [53]: {dmath} E_n, λ = (n + 12) + 34λ( 1 + 2n(n + 1)) - λ^2 ((n + 1) (n + 32) 2(n + 2)2 + 3 λ(2 n + 3) - n (n - 12) 2(n - 1)2 + 3 λ(2 n - 1) + (n + 1) (n + 2) (n + 3) (n + 4)16 (4 + 6 λ(2 n + 5)) - n (n - 1) (n - 2) (n - 3)16 (4 + 6 λ(2 n - 3))) As can be seen in table 3, the results agree for low $\lambda$ but diverge as the contribution of the quartic factor increases, which is exactly the behavior we expected.
A comparative plot of the obtained data is reported in Figure 3. First, we take a look at the dependence on $\lambda$ of the ground state and the state n = 2 (Figures 3a and 3b). The general shape of the states remains the same, however as $\lambda$ increases the state becomes more localized since the potential will grow more rapidly. Likewise, it is interesting to look at all the states for a given $\lambda$ (Figures 3c, 3d). Once again they have the same general shape as the harmonic states, as expected.
@@ V.2 Oscillator with imaginary frequency
For these given parameters, the energies of the two states are extremely close. As such, we had to impose a stricter convergence condition, stopping the training when the differential equation loss was less than $2\times 10^{-5}$. Regardless, the network did converge for all the studied states and we obtained an energy estimate of $-9.6808$ for the ground state and $-9.6806$ for the first excited state. As expected, these energies are negative as they correspond to bound states. Furthermore, by increasing $\lambda$ the eigenergies get closer to 0, as the potential becomes closer to a pure quartic oscillator and, therefore, the states get closer to unbound states. It is also useful to look at the shape of these eigenstates (Figure 4). These are, once again, what we expect from the shape of the potential: two peaks corresponding to the minima in the potential. The ground and the first excited state are also extremely similar, except for one symmetric and one anti-symmetric, and the peak with the same sign almost perfectly overlaps, thus justifying the closeness of the energy levels. Lastly, it is interesting to note, from Figure 5a that the local minima of the wavefunction corresponding to $x=0$ assumes a higher value as $\lambda$ increases. On the other hand, in the first excited we always have a node in $x=0$, see Figure 5b.
@@ VI Pure quartic oscillator
with $\lambda=1$. In order to initialize the network to a good state, we start with the weights and biases of the anharmonic oscillator with $\omega^{2}=1$ and $\lambda=1$. The network converged effectively for all states. In Figure 6 we show one of the excited states of the pure quartic oscillator compared with the state at the same quantum number for the harmonic oscillator. Once again the two have the same general shape, but the quartic one is much more localized since the potential grows faster with the position $x$.
@@ VII Energy behavior
The last result of our paper concerns the scaling behavior and the spacing of the energy levels in function of $\lambda$. It is indeed interesting to check whether our method is able to assess the departure from the validity of the perturbative regime. We compare in Figure 7 the eigenvalues of the anharmonic oscillator against the ones of the pure quartic oscillator. It is easy to see how the eigenvalues of the pure quartic oscillator increase at a roughly exponential rate when plotted against the logarithm of $\lambda$ (or, equivalently, they are proportional in log-log scale). On the other hand, we can recognize two regimes for the values of the anharmonic oscillator. Both seem to be roughly exponential, but for low values of $\lambda$ the energy increases much slower than for higher ones. Furthermore, for high $\lambda$ the anharmonic and pure quartic eigenvalues line up, since the harmonic contribution becomes negligible. We can get a rough idea of this behavior by performing a linear fit for the logarithms of $\lambda$ and E. In particular, we perform three fits: one for the low-$\lambda$ region (roughly $\lambda<0.1$) of the anharmonic oscillator, one for the high-$\lambda$ region (roughly $\lambda>2.0$) of the anharmonic oscillator, and lastly a fit for all the values of the quartic oscillator. We obtain, as expected, exponential behavior for each of the studied regions. Furthermore, we can see how the high-$\lambda$ behavior of the anharmonic oscillator matches the quartic oscillator for the last values of $\lambda$ when the quadratic part becomes negligible. See Appendix A for the numerical values of the fits.
@@
Considering the numerical analysis that we performed, a comment on the errors is in order. In our work there can be three major possible sources of uncertainty for the guess of the energy. The first is due to the stochasticity of the gradient descent. The second is due to the random initialization. We can estimate a lower bound for the magnitude of these two by running the training for the ground state of the harmonic oscillator ten times, and then looking at the standard deviation of the guesses for the energy. We obtain a $\sigma$ of 0.7% the value of the energy. The third is due to the choice of collocation points. Those are the ones that make up the uncertainty displayed in the results of Tables 2 and 3

### v2

@@ III.3 Choice of the domain
We need to choose an $L$ such that the wave function is vanishingly small at the boundaries, but at the same time, $L$ should be as small as possible to avoid training the network on areas of just zeros, which would make the network vastly underestimate the orthogonality loss (10) when using Mean Squared Error. This means that the optimal choice of $L$ depends on the state that we want to train the network to predict, since for this type of potential higher energy states will tend to be less localized and we must therefore choose a domain that grows as $n$ increases [^4] If this means that we need to extrapolate, for instance when obtaining the values of $\psi_{i}(x)$ for the orthogonality loss, we just predict 0 outside the original computational domain since the wave function should be negligible outside it.. In particular, we empirically find that a good value for $L$ for the unperturbed oscillator is:
@@
To that we add a constant $\epsilon=1$ on each side of the interval to ensure some leeway [^5] But if this makes $L_{a}(n,\lambda)>L(n)$, we just use $L(n)$., thus obtaining:
@@
Figure 1: Graphical representation of the approach followed in this paper. The circles represent variables, the rectangles represent operations. The variables are also represented by primary colors, while the methods are represented by the secondary colors obtained by combining the colors of their input and their output. In blue there are the neural network inputs. In purple there are the two neural networks, the main one that computes the eigenstates and the one that computes the energy. In red there are the outputs of the neural networks: the auxiliary outputs, the main outputs which is the eigenfunction and the eigenvalue $E$. In yellow there are the operations needed to calculate the losses. The losses are the integral loss, the boundary condition loss, the normalization loss, the inductive biases and the differential equation loss. SAE is the Sum of Absolute Errors, SSE is the Sum of Squared Errors. In green there is the final loss. It is calculated by summing the partial losses weighted by empirically adjusted hyperparameters $w_{int}$, $w_{bc}$, $w_{norm}$, $w_{bias}$ and $w_{eq}$. In cyan there are the criteria and the optimization method.
III.4 Materials
All computations have been performed on a MSI GF65 Thin 10UE laptop with a NVIDIA GeForce RTX 3060 Laptop GPU
@@ IV Results for the harmonic oscillator
We train the network to predict the first six stationary states. The training is stopped as soon as the total loss becomes less than $5\times 10^{-2}$ and the differential equation loss is smaller than $4\times 10^{-4}$, in order to obtain good guesses for the energies and eigenstates without requiring an overly long computational time. We set $E_{init}=0$ and $a=0.8$ in (12), and obtain the values reported below in Table 2:
@@
0 $3.76\times 10^{-4}$ $0.9999665$
1 $5.59\times 10^{-4}$ $0.9999776$
2 $-4.44\times 10^{-5}$ $0.9998280$
3 $1.18\times 10^{-5}$ $0.9999918$
4 $1.99\times 10^{-5}$ $0.9999005$
5 $5.10\times 10^{-5}$ $0.9999357$
@@
Figure 2: (a) Ground state for the harmonic oscillator. The red dots are the PINN’s predictions, while the blue line is the ground truth. (b) Fidelity throughout training for the ground state. (c) Loss behavior throughout training. The y axis is in logarithmic scale, therefore the oscillations on the y axis for low losses are overemphasized
@@
Figure 3: (a) Fifth excited state for the harmonic oscillator. The red dots are the PINN’s predictions, while the blue line is the ground truth. (b) Fidelity throughout training for the ground state. (c) Loss behavior throughout training. The y axis is in logarithmic scale, therefore the oscillations on the y axis for low losses are overemphasized
Figure 4: Behavior of the seven losses throughout training for the (a) ground state and (b) fifth excited state of the harmonic oscillator. The orthogonality loss is displayed in blue. Note that this loss cannot be seen in (a) as it is always 00 for the ground state. The symmetry loss is displayed in magenta. The integral loss is displayed in olive green. The normalization loss is displayed in orange. The boundary conditions loss is displayed in brown. The differential equation loss is displayed in black
@@
The error on the eigenenergy always stays beneath the 1%. The results for the fidelity are also more than acceptable, with a distance of at most in the order of $10^{-4}$, even though the fidelity slightly decreases as the energy increases. Some examples of the final results and the behavior of the loss and fidelity [^6] After each training step, the ground truth and the current PINN output are computed on a grid, and the fidelity between the two is computed. Note that this is a validation step that happens outside the training step, due to the unsupervised nature of the method. The network never actually utilizes the ground truth. throughout the training process are reported in Figure 2 for the ground state, and in Figure 3 for the fifth excited state. In particular, looking at the behavior of the loss (Figures 2c and 3c) we can see it is mostly monotonically decreasing, but for $n=5$ it takes roughly double the epochs then $n=0$ to get to the same value. Still, it shows good convergence. Concerning fidelity (Figures 2b and 3b), it displays excellent behavior for $n=5$, asymptotically approaching 1. On the other hand, the behavior for $n=0$ is more complex. It rapidly approaches $0.8$, then goes back down to almost zero, and eventually rises up until it reaches 1. A possible explanation could be that the network was approaching the ground state, then it risked converging to an excited state, which will be orthogonal to the ground and therefore with fidelity zero, but after several epochs it managed to converge to the correct state, thanks to the energy minimization condition. Furthermore, we can qualitatively see from Figures 2a and 3a that the neural network predictions seem to line up with the ground truth, as expected from the good values for the fidelity in Table 2. It is also interesting to look at the behavior of the six [^7] We exclude the energy minimization loss from this analysis, since it is only relevant in the initial epochs to avoid having the network converge to the wrong state. partial losses, in Figures 4. The overall behavior of the losses follows with good agreement the one of the total loss in Figures 2c and 3c. However, looking at the partial losses gives some important information. First of all, the differential equation and integral losses seem to oscillate far less compared to the other losses. This is something that must be taken into account when setting up the neural network’s parameters, especially the learning rate. Secondly, the losses relating to the physical constraints of the system seem to start stagnating after some epochs. This is not unexpected, since the heuristic detailed in Sec. III.1 sets up the weights in such a way that those conditions are satisfied before the network focuses on optimizing the differential equation.
Figure 5: Squared error relative to the position $x$ for the (a) ground state and (c) fifth excited state.
Lastly, we can also glean useful information by looking at the absolute error with respect to the position:
$Err_{\psi}(x)=|\psi_{Ex}(x)-\psi_{PINN}(x)|$ (20)
This error is displayed in Figure 5. In particular, for the ground state in Figure 5a we can see that the error tends to be mostly symmetric with respect to 00. This is probably a direct consequence both of the introduction of the symmetry loss and of the symmetric shape of the wavefunction itself. The same can be said for the fifth excited state in Figure 5b, with the difference that the scale of the error is much greater than that of the ground state. Time-wise, training a PINN takes $20$ to $40$ minutes for the ground state, and up to some hours for the higher excited states. This increase in computational time comes from both higher complexity for the shape of the excited states and the increase in the number of dot products that we need to compute in order to obtain the orthogonality loss.
@@ V.1 Oscillator with real frequency
We now add the quartic potential to the oscillator. In particular, we will use $\omega=1$ and train the neural network with $\lambda=0.005\times 2^{n}$, with $n$ from zero to $12$. These choices of $\lambda$ will allow us to study cases in which the quartic potential is just a perturbation, in which the two potentials have the same scales and in which the quartic part is dominant, up until the point where the system is basically just a pure quartic oscillator. In order to guarantee convergence, we exploit the results of section (IV) and employ transfer learning: we initialize the weight of the network to the ones that give the eigenfunction and eigenvalue of the harmonic oscillator, thus ensuring that the network starts close to the minima. Then, as $\lambda$ increases, we always initialize the network to the weight and biases as the already solved one with the closest $\lambda$. Due to this initialization, it becomes less important to enforce the orthogonality and energy minimization losses, since the minima closer to the starting state should already be the one corresponding to the correct eigenstate and eigenenergy, which means that we do not need these two losses as much to drive the network away from unwanted minima. Therefore, their starting weights have been decreased to $30n$ and 00 respectively. Furthermore, to ensure accurate results for this more complex system the convergence condition for the differential equation loss has been mede stricter, going from $4\times 10^{-4}$ to $2\times 10^{-5}.$ The network manages to reach the convergence conditions for all values of $\lambda$ and $n$. Thanks to transfer learning, these results were also reached much quicker. Time wise, it takes $5$ to $10$ minutes for lower excited states, compared to $20$ to $40$ minutes when given a clean initialization. Likewise, the convergence condition is reached much faster, with the harmonic oscillator requiring around $15000$ epochs starting from a clean initialization, while tha anharmonic oscillator with $\lambda=20.48$ takes only around $2000$ epochs.
@@
0 0.005 $2.06\times 10^{-2}$
5 0.005 $-8.67\times 10^{-4}$
2 0.16 $1.56\times 10^{-2}$
2 0.32 $-1.87\times 10^{-3}$
2 0.64 $-4.79\times 10^{-2}$
3 1.28 $-2.41\times 10^{-1}$
0 10.24 $-4.18\times 10^{-1}$
5 10.24 $-7.94\times 10^{-1}$
@@
At this point we compare the obtained eigenenergies with those one gets from the standard perturbation theory. In particular, we will have, for $\omega=1$ [64]: {dmath} E_n, λ = (n + 12) + 34λ( 1 + 2n(n + 1)) - λ^2 ((n + 1) (n + 32) 2(n + 2)2 + 3 λ(2 n + 3) - n (n - 12) 2(n - 1)2 + 3 λ(2 n - 1) + (n + 1) (n + 2) (n + 3) (n + 4)16 (4 + 6 λ(2 n + 5)) - n (n - 1) (n - 2) (n - 3)16 (4 + 6 λ(2 n - 3))) As can be seen in table 3, the results agree for low $\lambda$ but diverge as the contribution of the quartic factor increases, which is exactly the behavior we expected. Note that in this case the energy computed in perturbation theory does not constitute a ground truth, and the discrepancy of up to $79\%$ does not indicate poor performance of the network. This discrepancy comes from the fact that the assumptions of perturbation theory do not hold for such large values of $\lambda$. This leads to $E_{pert}$ to be largely incorrect while $E_{PINN}$ is probably much closer to the real value. To check the correctness of the results, we compare them to those obtained by the Legendre mesh method [65], which is known to display uncommonly high accuracy for the eigenergies of the anharmonic oscillator [66]. In particular, we utilize the implementation in the LagrangeMesh Mathematica package [67]. Some of the results are summarized in Table 4. The discrepancy remains below the $1\%$, confirming the efficacy of the model. It is important to note that utilizing Lagrange Mesh is both faster and more accurate then training a PINN. Utilization of PINNs should, however, show advantages for more complex systems, due to the expressive powers of neural networks. This analysis aims to be a building block towards these more complex applications. Another future application could be to train a network that takes as input both $x$ and $\lambda$, and outputs the $\psi(x)$ and $E$ for the anharmonic oscillator with that value of $\lambda$.
A comparative plot of the obtained data is reported in Figure 6. First, we take a look at the dependence on $\lambda$ of the ground state and the state $n=2$ (Figures 6a and 6b). The general shape of the states remains the same, however as $\lambda$ increases the state becomes more localized since the potential will grow more rapidly. Likewise, it is interesting to look at all the states for a given $\lambda$ (Figures 6c, 6d). Once again they have the same general shape as the harmonic states, as expected.
n $\lambda$ $err_{E}$ $\mathcal{F}_{\psi}$
0 0.005 $-5.78\times 10^{-4}$ $0.9999970$
5 0.005 $5.78\times 10^{-6}$ $0.9999962$
2 0.16 $2.29\times 10^{-3}$ $0.9999969$
2 0.32 $-6.28\times 10^{-4}$ $0.9999982$
2 0.64 $-8.25\times 10^{-5}$ $0.9999972$
3 1.28 $3.10\times 10^{-4}$ $0.99999921$
0 10.24 $-2.74\times 10^{-4}$ $0.999999974$
5 10.24 $-1.80\times 10^{-6}$ $0.9999999983$
Table 4: Results for the error on the energy between the state computed by the Lagrange Mesh method and the one computed neural network for some values of $n$ and $\lambda$ of the anharmonic oscillator.
@@ V.2 Oscillator with imaginary frequency
For these given parameters, the energies of the two states are extremely close. As such, we had to impose a stricter convergence condition compared to the harmonic oscillator, stopping the training when the differential equation loss was less than $2\times 10^{-5}$, as we did for the anharmonic oscillator with real frequency. The network converged for all the addressed states. It took after around $64000$ epochs for the ground state and around $112000$ epochs for the first excited state. We obtained an energy estimate of $-9.6808$ for the ground state and $-9.6806$ for the first excited state, the same values that can be obtained by the Lagrange Mesh method. As expected, these energies are negative as they correspond to bound states. Furthermore, by increasing $\lambda$ the eigenergies get closer to 0, as the potential becomes closer to a pure quartic oscillator and, therefore, the states get closer to unbound states. It is also useful to look at the shape of these eigenstates (Figure 7). These are, once again, what we expect from the shape of the potential: two peaks corresponding to the minima in the potential. The ground and the first excited state are also extremely similar, except for one symmetric and one anti-symmetric, and the peak with the same sign almost perfectly overlaps, thus justifying the closeness of the energy levels. The correctness of these eigenfunctions can also be confirmed by looking at their fidelity with the ones computed by the Lagrange Mesh method. The values of the fidelity are $0.99999984$ and $0.99999986$ for the ground state and first excited state, respectively. We also once again utilized transfer learning to train the network with different values of $\lambda$, and it is interesting to note, from Figure 8a that the local minima of the wavefunction corresponding to $x=0$ assumes a higher value as $\lambda$ increases. On the other hand, in the first excited we always have a node in $x=0$, see Figure 8b.
@@ VI Pure quartic oscillator
with $\lambda=1.28$. In order to initialize the network to a good state, we start with the weights and biases of the anharmonic oscillator with $\omega^{2}=1$ and $\lambda=1.28$. The network converged effectively for all states given the same convergence condition as the anharmonic oscillator. In Figure 9 we show one of the excited states of the pure quartic oscillator compared with the state at the same quantum number for the harmonic oscillator. Once again the two have the same general shape, but the quartic one is much more localized since the potential grows faster with the position $x$. We also once again check the correctness of the PINN’s predictions by comparing them to those obtained with the Lagrange mesh method, and verify their excellent agreement. These results are summarized in Table 5.
n $err_{E}$ $\mathcal{F}_{\psi}$
0 $3.57\times 10^{-4}$ 0.99999984
1 $5.39\times 10^{-4}$ 0.99999977
2 $3.03\times 10^{-6}$ 0.99999940
3 $-9.51\times 10^{-6}$ 0.99999986
4 $-1.06\times 10^{-5}$ 0.99999955
5 $-2.19\times 10^{-6}$ 0.999999986
Table 5: Results for the error on the energy and fidelity between the state computed by the Lagrange Mesh method and the one computed neural network for the pure quartic oscillator with $\lambda=1.28$.
@@ VII Energy behavior
The last result of our paper concerns the scaling behavior and the spacing of the energy levels as a function of $\lambda$. It is indeed interesting to check whether our method is able to assess the departure from the validity of the perturbative regime. We compare in Figure 10 the eigenvalues of the anharmonic oscillator against the ones of the pure quartic oscillator. It is easy to see how the eigenvalues of the pure quartic oscillator increase at a roughly exponential rate when plotted against the logarithm of $\lambda$ (or, equivalently, they are proportional in log-log scale). On the other hand, we can recognize two regimes for the values of the anharmonic oscillator. Both seem to be roughly exponential, but for low values of $\lambda$ the energy increases much slower than for higher ones. Furthermore, for high $\lambda$ the anharmonic and pure quartic eigenvalues line up, since the harmonic contribution becomes negligible. We can get a rough idea of this behavior by performing a linear fit for the logarithms of $\lambda$ and E. In particular, we perform three fits: one for the low-$\lambda$ region (roughly $\lambda<0.1$) of the anharmonic oscillator, one for the high-$\lambda$ region (roughly $\lambda>2.0$) of the anharmonic oscillator, and lastly a fit for all the values of the quartic oscillator. We obtain, as expected, exponential behavior for each of the studied regions. Furthermore, we can see how the high-$\lambda$ behavior of the anharmonic oscillator matches the quartic oscillator for the last values of $\lambda$ when the quadratic part becomes negligible. See Appendix A for the numerical values of the fits.